OpenStudy (anonymous):

I need help please. The formula E= 1000(100-T) + 580(100-T)^2 is used to approximate the elevation (in meters) above the sea level at which water boils at a temperature of T (in degrees Celsius). Find the rate of change of E with respect to T for a temperature of 51 Celsius.

OpenStudy (irishboy123):

\[ E= 1000(100-T) + 580(100-T)^2\] \[\frac{dE}{dT} = -1000 -1160(100-T)\] simplify that and plug in T = 51 it's been 11 hours so i am not going to apologise for feeding most of an answer but in future, maybe thing about using the word "calculus" in your opening post :p